# Gradient Networks

### v2

@@ 6.2 Cascaded (Monotone) Gradient Networks (GradNet-C, mGradNet-C)
The activation $\sigma_{\ell}$ in Thm. 8 can be a fixed function applied elementwise, e.g., tanh, softplus, and sigmoid, or a learnable function that operates on individual elements of the input vector. Thus, the proposed approach also permits elementwise activations parameterized by scalar-valued neural networks previously highlighted in Cor. 3.1. The proposed cascaded networks also remain GradNets if all $\boldsymbol{\beta}_{\ell}$ are 0. However, as shown in [55, 56], skip connections accelerate training as they address the vanishing gradient problem in deep networks and smoothen the loss landscape. Next we introduce the monotone variant of the GradNet-C.
@@ 7.1 Gradient Field
Contour plots of $F$ and $G$, along with their corresponding gradients, are shown in Fig. 4. For the convex test function $F(\boldsymbol{x})$ in (38), we compare the proposed mGradNet-C and mGradNet-M, respectively defined in (33)-(35) and (30)-(31), with the Input Convex Neural Network (ICNN) [36], Input Convex Gradient Network (ICGN) [29], and Convex Ridge Regularizer (CRR) [34]. Similarly for the nonconvex test function $G(\boldsymbol{x})$ in (39), we compare the proposed modular gradient network (GradNet-M) and cascaded gradient network (GradNet-C) to a multilayer perceptron (MLP) and a ridge regularizer (RR) network proposed in [35]. The input to all models is a point $\boldsymbol{x}\in[0,1]^{2}$ and all networks, excluding the MLP and ICNN, are trained to directly output $\nabla F(\boldsymbol{x})$ or $\nabla G(\boldsymbol{x})$. The input to the ICNN and MLP is also $\boldsymbol{x}$ and the gradient of the networks, with respect to $\boldsymbol{x}$, are trained to approximate $\nabla F(\boldsymbol{x})$ or $\nabla G(\boldsymbol{x})$ respectively. The specific architecture configurations for each model are provided in Appendix A. All models are constrained to have roughly 100 parameters. We sample 100,000 training points uniformly at random on the unit square and separately sample 10,000 validation points. All models are trained using mean squared error (MSE) loss and their performance is evaluated on a uniform grid of points in the unit square. The $\ell_{2}$ norm of the gradient prediction error is shown for each model in Fig. 5.
@@
We observe from Fig. 5 that, for learning $\nabla F(\boldsymbol{x})$, the proposed mGradNet-M and mGradNet-C exhibit significantly lower error rates than that of the ICNN, ICGN, and CRR. Fig. 5 also shows that the mGradNet-M and mGradNet-C are able to effectively learn the gradient at all areas of the unit square, whereas the ICNN, ICGN, and CRR tend to underperform around the edges of square. Similar trends are observed in the nonconvex case where the GradNet-C, GradNet-M, and RR [35] outperform the MLP.
@@ 7.1.2 $d$-Dimensional Gradient Field, $d\geq 32$
In (44) above, $N(\boldsymbol{x};\boldsymbol{\mu}_{i};\boldsymbol{\Sigma}_{i})$ denotes the probability density function of a multivariate normal distribution with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{d}$ and $d\times d$ covariance matrix $\boldsymbol{\Sigma}_{i}$. Since most of the probability mass of a $d$-dimensional Gaussian lies in a thin annulus that is centered at the origin and has inner and outer radii close to $\sigma\sqrt{d}$ [58], we ensure interesting score functions in the unit hypercube by drawing each component of $\boldsymbol{\mu}_{i}$ uniformly at random from $[0.3,0.7]$ and setting $\boldsymbol{\Sigma}=2\sqrt{d}\boldsymbol{I}$. We select equal weights for the Gaussians. We train all models in these experiments for 10,000 iterations, where at each iteration we randomly sample 100 points from $[0,1]^{d}$ to update the model parameters. The specific model configurations and additional training details are provided in Appendix B.
@@
Similarly for the the score learning task, we observe that in Table 2 that the GradNet-M and GradNet-C significantly outperform all other methods. The GradNet-C outperforms the RR by nearly 8 dB for $d=32$ and over 4 dB for $d=256$. The GradNet-M outperforms the RR by a margin of over 5 dB for $d=1024$ and demonstrates its effectiveness in high dimensions. The MLP baseline underperformed all other considered methods, especially for $d=1024$, where its performance lagged by nearly 9 dB behind the second worst model, namely the RR. In both the convex and nonconvex experiments, the modular architectures, mGradNet-M and GradNet-M, perform better than the other methods in the high-dimensional setting $d=1024$.
@@ 7.2 Hamiltonian Gradients for Two-Body Dynamics
In this section, we train GradNets to learn gradients of the Hamiltonian function for the two-body problem and subsequently use the learned models to predict the trajectories of the objects. We observe that our GradNet-M outperforms the baseline Hamiltonian NN [59] by over 3 dB and the RR [35] by over 11 dB. As in the gradient field experiments, our GradNet-C architecture again outperforms the RR [35] – this time by over 3 dB. The performance gaps that arise in this experiment directly validate the theoretical differences in approximation capabilities between the baseline methods and GradNets (see Sec. 5).
@@
In Fig. 6, we plot the ground truth trajectories of the two bodies along with the trajectories unrolled using the estimated gradients of the Hamiltonian in (46). Table 3 shows the corresponding coordinate MSEs for the unrolled trajectories and reflects the results in Fig. 6. Since the energy in the system should remain constant, the energy MSEs in Table 3 reflect how effectively the Hamiltonian gradient predictions of each method conserve the total energy of the system. We see that the GradNet-M outperforms the Hamiltonian NN baseline by over 3 dB in terms of coordinate MSE and by approximately 1 dB in terms of energy MSE. Similarly, the GradNet-M outperforms the RR by 11 dB for coordinate MSE and 32 dB for energy MSE. We attribute this gain to the guaranteed universal approximation capabilities of of the GradNet-M discussed in Sec. 5. The GradNet-C outperforms the RR and demonstrates better approximation capabilities for the gradient of (46), which is not a sum of ridge functions.

### v3

@@ 6.2 Cascaded (Monotone) Gradient Networks (GradNet-C, mGradNet-C)
The activation $\sigma_{\ell}$ in Thm. 8 can be a fixed function applied elementwise, e.g., tanh, softplus, and sigmoid, or a learnable function that operates on individual elements of the input vector. Thus, the proposed approach also permits elementwise activations parameterized by scalar-valued neural networks previously highlighted in Cor. 3.1. The proposed cascaded networks also remain GradNets if all $\boldsymbol{\beta}_{\ell}$ are 0. However, as shown in [55, 56], skip connections accelerate training as they address the vanishing gradient problem in deep networks and smoothen the loss landscape. Next, we introduce the monotone variant of the GradNet-C.
@@ 7.1 Gradient Field
Contour plots of $F$ and $G$, along with their corresponding gradients, are shown in Fig. 4. For the convex test function $F(\boldsymbol{x})$ in (38), we compare the proposed mGradNet-C and mGradNet-M, respectively defined in (33)-(35) and (30)-(31), with the Input Convex Neural Network (ICNN) [36], Input Convex Gradient Network (ICGN) [29], and Convex Ridge Regularizer (CRR) [34]. Similarly for the nonconvex test function $G(\boldsymbol{x})$ in (39), we compare the proposed modular gradient network (GradNet-M) and cascaded gradient network (GradNet-C) to a multilayer perceptron (MLP) and the ridge regularizer (RR) network proposed in [35]. The input to all models is a point $\boldsymbol{x}\in[0,1]^{2}$. All networks, excluding the ICNN and MLP, are trained to directly output $\nabla F(\boldsymbol{x})$ or $\nabla G(\boldsymbol{x})$. The input to the ICNN and MLP is also $\boldsymbol{x}$ and the gradient of the networks, with respect to $\boldsymbol{x}$, are trained to approximate $\nabla F(\boldsymbol{x})$ or $\nabla G(\boldsymbol{x})$, respectively. The specific architecture configurations for each model are provided in Appendix A. All models are constrained to have roughly 100 parameters. We sample 100,000 training points uniformly at random on the unit square and separately sample 10,000 validation points. All models are trained using mean squared error (MSE) loss and their performance is evaluated on a uniform grid of points in the unit square. The $\ell_{2}$ norm of the gradient prediction error is shown for each model in Fig. 5.
@@
We observe from Fig. 5 that, for learning $\nabla F(\boldsymbol{x})$, the proposed mGradNet-M and mGradNet-C exhibit significantly lower error rates than that of the ICNN, ICGN, and CRR. Fig. 5 also shows that the mGradNet-M and mGradNet-C are able to effectively learn the gradient at all areas of the unit square, whereas the ICNN, ICGN, and CRR tend to underperform around the edges of square. Similar trends are observed in the nonconvex case where the GradNet-C, GradNet-M, and RR outperform the MLP.
@@ 7.1.2 $d$-Dimensional Gradient Field, $d\geq 32$
In (44) above, $N(\boldsymbol{x};\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$ denotes the probability density function of a multivariate normal distribution with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{d}$ and $d\times d$ covariance matrix $\boldsymbol{\Sigma}_{i}$. Since most of the probability mass of a $d$-dimensional standard normal distribution lies in a thin annulus that is centered at the origin and has inner and outer radii close to $\sqrt{d}$ [58], we ensure interesting score functions in the unit hypercube by drawing each component of $\boldsymbol{\mu}_{i}$ uniformly at random from $[0.3,0.7]$ and setting $\boldsymbol{\Sigma}=2\sqrt{d}\boldsymbol{I}$. We select equal weights for the Gaussians. We train all models in these experiments for 10,000 iterations, where at each iteration we randomly sample 100 points from $[0,1]^{d}$ to update the model parameters. The specific model configurations and additional training details are provided in Appendix B.
@@
Similarly for the the score learning task, we observe in Table 2 that the GradNet-M and GradNet-C significantly outperform all other methods. The GradNet-C outperforms the RR by nearly 8 dB for $d=32$ and by over 4 dB for $d=256$. The GradNet-M outperforms the RR by a margin of over 5 dB for $d=1024$, thus demonstrating its effectiveness in high dimensions. The MLP baseline underperformed all other considered methods, especially for $d=1024$, where its performance lagged by nearly 9 dB behind the second worst model, namely the RR. In both the convex and nonconvex experiments, the modular architectures, mGradNet-M and GradNet-M, perform better than the other methods in the high-dimensional setting $d=1024$.
@@ 7.2 Hamiltonian Gradients for Two-Body Dynamics
In this section, we train GradNets to learn gradients of the Hamiltonian function for the two-body problem and subsequently use the learned models to predict the trajectories of the two objects. We observe that our GradNet-M outperforms the baseline Hamiltonian NN [59] by over 3 dB and the RR [35] by over 11 dB. As in the gradient field experiments, our GradNet-C architecture again outperforms the RR [35] – this time by over 3 dB. The performance gaps that arise in this experiment directly validate the theoretical differences in approximation capabilities between the baseline methods and GradNets (see Sec. 5).
@@
In Fig. 6, we plot the ground truth trajectories of the two bodies along with the trajectories unrolled using the estimated gradients of the Hamiltonian in (46). Table 3 shows the corresponding coordinate MSEs for the unrolled trajectories and reflects the results in Fig. 6. Since the energy in the system should remain constant, the energy MSEs in Table 3 reflect how effectively the Hamiltonian gradient predictions of each method conserve the total energy of the system. We see that the GradNet-M outperforms the Hamiltonian NN baseline by over 3 dB in terms of coordinate MSE and by approximately 1 dB in terms of energy MSE. Similarly, the GradNet-M outperforms the RR by 11 dB for coordinate MSE and 32 dB for energy MSE. We attribute this gain to the universal approximation capabilities of of the GradNet-M discussed in Sec. 5. The GradNet-C outperforms the RR and demonstrates better approximation capabilities for the gradient of (46), which is not a sum of ridge functions.